the average human being contains 60 trillion living cells. if there are currently 7.217 million people on earth. how many moles of living human cells are currently on our planet? if 150billion humans have inhabited the planet in all of history, how many moles of human cells have existed so far on our planet?
1
Expert's answer
2020-07-22T05:41:27-0400

According to the definition of the mole, one mole of the living cells contains the Avogadro number of the living cells "N_A = 6.022\u00b710^{23}" mol-1. Therefore, the number of the moles of living human cells currently on our planet is:

"n = \frac{N}{N_A} = \frac{60·10^{18}·7.217·10^6}{6.022·10^{23}} = 719" mol.

If 150 billion humans have inhabited the planet in all of history, the number of the moles of human cells have existed so far on our planet is:

"n_1 = \frac{N_1}{N_A} = \frac{60·10^{18}·150·10^9}{6.022·10^{23}} = 1.49·10^7" mol.